### 16 Apr 2026 v1

# Boson star-black hole binaries: initial data and head-on collisions

Abstract
We present a numerical-relativity study of comparable-mass boson star-black hole (BS-BH) head-on collisions, focusing on both initial-data construction and gravitational-wave (GW) phenomenology. We show that plain superposition can strongly perturb the BS core, leading to large constraint violations and unphysical radial oscillations. To remedy this problem, we introduce a one-body conformal-factor correction and find that it robustly suppresses these artifacts. Using the improved initial data, we analyze GW emission from equal- and unequal-mass BS-BH binaries and compare with matched BS-BS and BH-BH baselines. For equal masses, the BS-BH radiated energy increases with BS compactness and approaches the BH-BH limit for highly compact stars. For unequal masses, the dominant $(2,0)$ mode often remains close to the BH-BH morphology, whereas the subdominant $(3,0)$ mode provides clear discriminatory power when the BH is the heavier companion. Our results identify higher multipoles as a key observable for distinguishing mixed BS-BH mergers from pure BH binaries.
I Introduction
The existence of dark matter (DM) is supported by a broad range of gravitational observations Navarro et al. (1996); Clowe et al. (2006), while its microphysical nature remains unknown and constitutes one of the most important open problems in modern cosmology and fundamental physics Bertone et al. (2005); Feng (2010). Despite extensive laboratory and astrophysical searches, no conclusive non-gravitational detection channel has yet been established Feng (2010); Klasen et al. (2015); Kahlhoefer (2017); Cebrián (2023). In this context, gravitational-wave (GW) astronomy Abbott and others (2019, 2023); Abac and others (2025) offers a complementary and potentially unique avenue Brito et al. (2017); Bertone and others (2020): if some fraction of DM can form compact self-gravitating structures, their mergers may leave directly observable imprints in the strong-gravity regime.
Several DM candidates have been proposed, ranging from particle-like candidates such as weakly interacting massive particles (WIMPs) Pospelov et al. (2008); Arcadi et al. (2018); Roszkowski et al. (2018); Schumann (2019) and sterile neutrinos Dodelson and Widrow (1994); Shi and Fuller (1999); Abazajian and others (2012), to wave-like candidates such as ultralight scalar fields Hu et al. (2000); Hui et al. (2017); Hui (2021), and even macroscopic objects such as primordial black holes (PBHs) Carr et al. (2016); Green and Kavanagh (2021); Ning et al. (2026b); Zeng et al. (2025); Carr et al. (2026); Ning et al. (2026a); Klipfel et al. (2026); Jia et al. (2026). Among these candidates, scalar fields are particularly well motivated and have been studied extensively. They arise naturally in a wide range of theoretical scenarios, including the QCD axion Dine et al. (1981); Dine and Fischler (1983); Preskill et al. (1983), axion-like particles Marsh (2016) from string-theory compactifications Svrcek and Witten (2006); Arvanitaki et al. (2010), and more general scalar-field DM models. If endowed with mass, scalar fields minimally coupled to gravity can form macroscopic self-gravitating structures Kaup (1968); Ruffini and Bonazzola (1969); Seidel and Suen (1991). Complex scalar fields can support static, spherically symmetric configurations in which the field oscillates in time, known as boson stars (BSs) Kaup (1968); Ruffini and Bonazzola (1969), while real scalar fields can instead form long-lived oscillating configurations with a nontrivial time-dependent energy-momentum tensor, called oscillatons Seidel and Suen (1991) (see, e.g., Refs. Schunck and Mielke (2003); Liebling and Palenzuela (2023); Visinelli (2021); Bezares and Sanchis-Gual (2025) for reviews). Both types of solutions can arise naturally as end states of gravitational collapse Kaup (1968); Ruffini and Bonazzola (1969); Seidel and Suen (1991); Okawa et al. (2014a); Chavanis (2016); Widdicombe et al. (2018) and have a maximum mass scaling as $M_{\mathrm{max}}\sim M_{\mathrm{pl}}^{2}/\mu$, where $M_{\mathrm{pl}}$ is the Planck mass and $\mu$ is the scalar-field mass. Consequently, depending on the scalar mass, these objects can span a wide range of masses and play very different cosmological roles.
Ultralight scalar fields with masses in the range $\mu\sim 10^{-22}-10^{-20}\,\mathrm{eV}$, often referred to as “fuzzy” DM, have de Broglie wavelengths on astrophysical scales and can therefore modify halo structure on kiloparsec scales, potentially alleviating some of the small-scale tensions of collisionless cold-DM phenomenology Hu et al. (2000); Hui et al. (2017); Hui (2021). The corresponding solitonic halo cores Schive et al. (2014); Veltmaat et al. (2018) are mathematically related to self-gravitating bosonic configurations, but they are typically much less compact and much larger than relativistic compact stars.
At the same time, axion and axion-like scenarios, as well as string-inspired “axiverse” constructions, motivate a much broader mass range of scalar fields. For heavier scalars, such as $\mu\sim 10^{-10}-10^{-16}\,\mathrm{eV}$, the associated self-gravitating configurations can span masses from solar to supermassive scales. In addition, modifying the scalar self-interaction potential can further enlarge the allowed ranges of mass and compactness, enabling configurations that behave as exotic compact objects (ECOs) or, in some regimes, as black-hole (BH) mimickers Cardoso et al. (2014); Marks et al. (2025); Staelens (2026); Evstafyeva et al. (2026). The mergers of such objects can produce strong GW signals that may be detectable by current and future GW observatories, and these signals may be distinguishable from standard BH or neutron-star (NS) mergers through their characteristic features.
Motivated by these properties, BS-BS binary mergers have been extensively investigated in numerical relativity, including head-on collisions Bernal and Siddhartha Guzman (2006); Palenzuela et al. (2007); Choptuik and Pretorius (2010); Mundim (2010); Cardoso et al. (2016); Helfer et al. (2022); Jaramillo et al. (2022); Sanchis-Gual et al. (2022b); Evstafyeva et al. (2023b); Atteneder et al. (2024); Siemonsen (2024); Lira et al. (2025); Ge et al. (2025); Brito et al. (2026); Ge (2025) and orbital inspirals Palenzuela et al. (2008); Mundim (2010); Palenzuela et al. (2017); Bezares et al. (2017); Bezares and Palenzuela (2018); Sanchis-Gual et al. (2020); Bezares et al. (2022); Croft et al. (2023); Siemonsen and East (2023a, b) for both equal- Palenzuela et al. (2007, 2008); Choptuik and Pretorius (2010); Mundim (2010); Cardoso et al. (2016); Palenzuela et al. (2017); Bezares et al. (2017); Bezares and Palenzuela (2018); Sanchis-Gual et al. (2020); Helfer et al. (2022); Jaramillo et al. (2022); Sanchis-Gual et al. (2022b); Croft et al. (2023); Siemonsen and East (2023a); Atteneder et al. (2024); Lira et al. (2025); Ge et al. (2025); Brito et al. (2026); Ge (2025) and unequal-mass Bernal and Siddhartha Guzman (2006); Bezares et al. (2022); Evstafyeva et al. (2023b); Siemonsen and East (2023b); Siemonsen (2024) binaries. These studies have revealed rich phenomenology, including waveform degeneracies with BH binaries and the imprint of tidal deformability. Furthermore, other ECO mergers have also been studied, such as oscillaton Helfer et al. (2019); Widdicombe et al. (2020) and Proca-star (vector cousins of BSs) Sanchis-Gual et al. (2019); Calderón Bustillo et al. (2021); Calderon Bustillo et al. (2023); Sanchis-Gual et al. (2022a) binaries. From an astrophysical perspective, if BSs, fuzzy-DM cores, or other scalar-field configuratons populate the Universe, their coexistence with BHs and their inevitable interactions are highly probable. In this context, the nonlinear dynamics of scalar fields coupled to BHs—including scalar-field accretion Sanchis-Gual et al. (2015); Hui et al. (2019); Clough et al. (2019); Cardoso et al. (2022a); Brito et al. (2015a, 2016), scalarization Okawa et al. (2014b); Silva et al. (2018); Zhang et al. (2022); Doneva et al. (2024); Chen et al. (2023b); Guo et al. (2026), descalarization Silva et al. (2021); Chen et al. (2023a), superradiance Brito et al. (2015b); Baryakhtar et al. (2021); Lyu et al. (2025), dynamical friction Traykova et al. (2021); Boudon et al. (2022), and dynamical transition processes Ning et al. (2024); Chen et al. (2023c); Zhao et al. (2026b, a)—have been widely explored in various strongly gravitating systems, revealing rich time-evolution behaviors and phase structures. It is therefore also crucial to understand the dynamics and GW signatures of mixed BS-BH binary mergers for interpreting future observations and constraining the existence of such objects.
Compared with BS-BS binaries, BS-BH binaries remain significantly less explored. Ref. Clough et al. (2018) presented the first fully relativistic simulations of head-on collisions between oscillatons (axion stars, ASs) and BHs/NSs, finding that in the AS-BH case the remaining scalar cloud after merger can be enhanced by decreasing the compactness of the AS and increasing the BH spin. For BS-BH binaries, Refs. Cardoso et al. (2022b); Zhong et al. (2023) studied piercing collisions, focusing on the large-mass-ratio regime, in which a small BH pierces a much larger BS. They found that the scalar field is almost entirely accreted by the BH for both mini and solitonic BSs. In addition, a small fraction of the scalar field survives outside the BH horizon and forms a long-lived quasi-bound state comoving with the BH, i.e., a “gravitational atom”.
However, as discussed above, BS masses can span a wide range, and it is plausible that they may be comparable to those of BHs. In this regime, the dynamics of BS-BH mergers can differ significantly from the large-mass-ratio case and may produce distinct GW signatures. In this work, we focus on head-on collisions of comparable-mass BS-BH binaries and study how the mass ratio and binary type (BS-BH versus BS-BS and BH-BH) affect GW emission. Very recently, during the final stage of this work, Ref. Marks et al. (2026) appeared. The authors also studied comparable-mass BS-BH collisions, examining GW radiation efficiency across different scalar potentials, BS compactnesses, velocities, and mass ratios. They also presented the first simulation of BS-BH inspirals, showing that an appropriate scalar self-interaction can suppress tidal disruption. In this work, we use a different initial-data construction method and different diagnostics to analyze the merger dynamics and GW signatures. Our results are complementary to those of Ref. Marks et al. (2026), and together they provide a more comprehensive understanding of comparable-mass BS-BH mergers.
From a numerical standpoint, the comparable-mass regime is also one in which the quality of the initial data becomes critical. A direct pointwise superposition of boosted single-object solutions (plain superposition) is simple and widely used Cardoso et al. (2022b); Zhong et al. (2023), but for BSs it can induce unphysical distortions of the stellar core volume element, increase constraint violations, and trigger spurious premature collapse. This issue has already been identified and cured for BS-BS binaries through metric correction and conformal-factor correction in the equal- Helfer et al. (2022) and unequal-mass Evstafyeva et al. (2023b) cases, respectively, leading to significant improvements in initial-data quality. The problem is expected to be even more severe for BS-BH binaries, since the gravitational tail of a BH is much stronger than that of a BS.
In Ref. Marks et al. (2026), the authors mitigated this issue by employing a metric correction method for the BS and a TwoPunctures correction for the BH. In this work, developed independently, we adopt a BS-centered one-body conformal-factor correction derived directly from the unequal-mass BS-BS prescription in the appropriate BS-BH limit. This method is self-consistent and provides a unified framework for constructing initial data for BS-BH and BS-BS binaries on an equal footing, allowing for a more direct comparison of their dynamics and GW signatures. We show that our correction strongly suppresses the initial constraint violations near the BS core and eliminates the spurious premature collapse observed with plain superposition. We then use these improved data to study the merger dynamics and GW emission of BS-BH binaries and compare their behavior with that of BS-BS and BH-BH counterparts.
The paper is organized as follows. In Sec. II, we summarize the Einstein-Klein-Gordon model, the $3+1$ formulation, and our numerical infrastructure and diagnostics. Section III presents isolated objects, boost construction, plain superposition, and the improved BS-BH prescription. In Sec. IV, we analyze the quality of the initial data and study equal- and unequal-mass head-on collisions. Finally, we conclude in Sec. V. Throughout this work, we use natural units $\hbar=c=G=1$ and code units with $\mu=1$ unless stated otherwise. Spacetime indices are denoted by Greek letters ($0\ldots 3$), and spatial indices by Latin letters ($1\ldots 3$).
II Formalism
II.1 Model
The action for a self-gravitating complex scalar field $\varphi$ minimally coupled to gravity is given by
$S=\int\mathrm{d}^{4}x\sqrt{-g}\left[\frac{R}{16\pi}-\frac{1}{2}\left(g^{\mu\nu}\nabla_{\mu}\bar{\varphi}\nabla_{\nu}\varphi+V(|\varphi|^{2})\right)\right],$ (1)
where $g$ is the determinant of the spacetime metric $g_{\mu\nu}$, $R$ and $\nabla_{\mu}$ denotes the Ricci scalar and the covariant derivative associated with this metric, $\bar{\varphi}$ is the complex conjugate of the scalar field, and $V(|\varphi|^{2})$ is the scalar-field potential. Varying the action with respect to the metric and the scalar field yields the Einstein-Klein-Gordon equations:
$R_{\mu\nu}-\frac{1}{2}Rg_{\mu\nu}$ $=8\pi T_{\mu\nu},$ (2)
$\nabla^{\mu}\nabla_{\mu}\varphi$ $=V^{\prime}\varphi\equiv\frac{dV}{d|\varphi|^{2}}\varphi,$ (3)
where the energy-momentum tensor of the complex scalar field is
$T_{\mu\nu}$ $=\frac{1}{2}\left(\nabla_{\mu}\bar{\varphi}\nabla_{\nu}\varphi+\nabla_{\nu}\bar{\varphi}\nabla_{\mu}\varphi\right)$ (4)
$\quad-\frac{1}{2}g_{\mu\nu}\left(g^{\alpha\beta}\nabla_{\alpha}\bar{\varphi}\nabla_{\beta}\varphi+V(|\varphi|^{2})\right).$
BSs are stationary solutions to the Einstein-Klein-Gordon equations, and their properties and compactness depend crucially on the choice of the potential $V(|\varphi|^{2})$. In this work, we focus on the solitonic potential Lee and Pang (1987):
$V=\mu^{2}|\varphi|^{2}\left(1-2\frac{|\varphi|^{2}}{\sigma_{0}^{2}}\right)^{2},$ (5)
where $\mu$ denotes the bare mass of the scalar field and $\sigma_{0}$ regulates the strength of the self-interaction. Note that the non-self-interacting mini BS is recovered in the limit $\sigma_{0}\to\infty$. In our numerical implementation, we adopt code units with $\mu=1$, so all time and length variables are expressed in units of $\mu^{-1}$.
II.2 $3+1$ formulation and computational infrastructure
We evolve the Einstein-Klein-Gordon equations using the standard $3+1$ decomposition of spacetime Gourgoulhon (2007); Alcubierre (2008); Shibata (2016), which decomposes the spacetime into spatial metric $\gamma_{ij}$, lapse function $\alpha$, and shift vector $\beta^{i}$ in adapted coordinates $(t,x^{i})$. The line element can be written as
$ds^{2}=-\alpha^{2}dt^{2}+\gamma_{ij}(dx^{i}+\beta^{i}dt)(dx^{j}+\beta^{j}dt).$ (6)
To cast the Einstein-Klein-Gordon equations into a first-order-in-time system, we introduce the extrinsic curvature
$K_{ij}\equiv-\frac{1}{2\alpha}(\partial_{t}\gamma_{ij}-\mathcal{L}_{\beta}\gamma_{ij})=-\frac{1}{2\alpha}(\partial_{t}\gamma_{ij}-D_{i}\beta_{j}-D_{j}\beta_{i}),$ (7)
and the conjugate momentum of the scalar field
$\Pi\equiv-\frac{1}{\alpha}(\partial_{t}\varphi-\mathcal{L}_{\beta}\varphi)=-\frac{1}{\alpha}(\partial_{t}\varphi-\beta^{i}\partial_{i}\varphi),$ (8)
where $\mathcal{L}_{\beta}$ denotes the Lie derivative along the shift vector $\beta^{i}$, and $D_{i}$ is the covariant derivative compatible with the spatial metric $\gamma_{ij}$. The evolution equations for $\gamma_{ij}$, $K_{ij}$, $\varphi$, and $\Pi$ then follow from the Einstein-Klein-Gordon equations:
$\partial_{t}\gamma_{ij}$ $=-2\alpha K_{ij}+\mathcal{L}_{\beta}\gamma_{ij},$ (9a)
$\partial_{t}K_{ij}$ $=-D_{i}D_{j}\alpha+\alpha\left(\mathcal{R}_{ij}+KK_{ij}-2K_{ik}{K^{k}}_{j}\right$.
$\left.\quad+4\pi\left[(S-\rho)\gamma_{ij}-2S_{ij}\right]\right)+\mathcal{L}_{\beta}K_{ij},$ (9b)
$\partial_{t}\varphi$ $=-\alpha\Pi+\mathcal{L}_{\beta}\varphi,$ (9c)
$\partial_{t}\Pi$ $=\alpha\left(K\Pi+V^{\prime}\varphi-D_{i}D^{i}\varphi\right)$
$\quad-(D_{i}\alpha)D^{i}\varphi+\mathcal{L}_{\beta}\Pi.$ (9d)
These equations are subject to the Hamiltonian and momentum constraints:
$\mathcal{H}$ $\equiv\mathcal{R}+K^{2}-K_{ij}K^{ij}-16\pi\rho=0,$ (10)
$\mathcal{M}_{i}$ $\equiv D_{j}{K^{j}}_{i}-D_{i}K-8\pi j_{i}=0,$ (11)
where $K\equiv\gamma^{ij}K_{ij}$, and $\mathcal{R}_{ij}$ and $\mathcal{R}$ denote the Ricci tensor and Ricci scalar associated with the spatial metric $\gamma_{ij}$, respectively. The source terms $\rho$, $j_{i}$, $S_{ij}$, and $S$ are defined as
$\rho$ $\equiv n_{a}n_{b}T^{ab}=\frac{1}{2}(\bar{\Pi}\Pi+\partial^{a}\bar{\varphi}\partial_{a}\varphi+V),$ (12)
$j_{i}$ $\equiv-\gamma_{ia}n_{b}T^{ab}=\frac{1}{2}(\bar{\Pi}\partial_{i}\varphi+\Pi\partial_{i}\bar{\varphi}),$
$S_{ij}$ $\equiv\gamma_{ia}\gamma_{jb}T^{ab}=\frac{1}{2}(\partial_{i}\bar{\varphi}\partial_{j}\varphi+\partial_{j}\bar{\varphi}\partial_{i}\varphi)$
$\quad-\frac{1}{2}\gamma_{ij}\left(\partial^{a}\bar{\varphi}\partial_{a}\varphi-\bar{\Pi}\Pi+V\right),$
$S$ $\equiv\gamma^{ij}S_{ij}=\frac{1}{2}(3\bar{\Pi}\Pi-\partial^{a}\bar{\varphi}\partial_{a}\varphi-3V).$
For stable numerical evolutions, the evolution system must be written in a strongly hyperbolic form. To this end, we use the publicly available code GRChombo Clough et al. (2015); Radia et al. (2022); Andrade and others (2021), which is built on the Chombo framework for adaptive mesh refinement (AMR) Adams and others (2015) and evolves the Einstein equations using the Baumgarte-Shapiro-Shibata-Nakamura-Oohara-Kojima (BSSNOK) Nakamura et al. (1987); Shibata and Nakamura (1995); Baumgarte and Shapiro (1998) and covariant conformal Z4 (CCZ4) Alic et al. (2012, 2013) formulations. In this work, we choose the CCZ4 formulation because of its constraint-damping properties. Specifically, we describe the spacetime in terms of conformally rescaled and trace-split variables,
$\chi$ $\equiv(\det\gamma_{ij})^{-1/3},$ (13)
$\tilde{\gamma}_{ij}$ $\equiv\chi\gamma_{ij},$
$\tilde{A}_{ij}$ $\equiv\chi\left(K_{ij}-\frac{1}{3}\gamma_{ij}K\right),$
$\tilde{\Gamma}^{i}$ $\equiv\tilde{\gamma}^{ab}\tilde{\Gamma}^{i}_{ab},$
and we choose $\kappa_{1}\rightarrow\kappa_{1}/\alpha$, $\kappa_{1}=0.1$, $\kappa_{2}=0$, and $\kappa_{3}=1$ for the constraint-damping parameters Radia et al. (2022). The full Einstein equations in the CCZ4 formulation can be found in Sec. III F of Ref. Alic et al. (2013). In addition, we use the ExoZvezda Helfer et al. (2022); Croft (2023); Croft et al. (2023); Evstafyeva et al. (2023b, a, 2024) code to compute one-dimensional BS solutions in isotropic gauge and to dynamically evolve the complex scalar field.
GRChombo discretizes the CCZ4 equations using the method of lines, with fourth-order finite-difference stencils in space and a fourth-order Runge-Kutta integrator in time. Unless otherwise indicated, we choose the Courant-Friedrichs-Lewy (CFL) factor to be $0.25$ and the Kreiss-Oliger dissipation coefficient to be $0.3$. We use the second derivatives of the complex scalar field $\varphi$ and the conformal factor $\chi$ as tagging criteria for AMR, with thresholds of $0.5$ and $0.25$, respectively. We employ a domain of length $512$ with a nested grid hierarchy of $L=7$ refinement levels, with grid spacing $dx=1/32$ on the finest level and a factor of $2$ increase between successive levels farther out. For GW extraction, we choose an extraction radius of $R_{\mathrm{ex}}=120$. We impose reflective boundary conditions on the $y=0$ and $z=0$ planes, since the collision always occurs along the $x$ axis, and radiative (outgoing) boundary conditions on the remaining boundaries. We use the standard $1+\mathrm{log}$ and Gamma-driver gauge conditions Alcubierre (2008). Convergence tests are presented in Appendix A.
II.3 Diagnostics
The $U(1)$ symmetry of the complex scalar-field action gives rise to a conserved current and an associated Noether charge,
$J^{\mu}$ $=\frac{i}{2}(\varphi\nabla^{\mu}\bar{\varphi}-\bar{\varphi}\nabla^{\mu}\varphi),$ (14)
$N$ $=\int\mathrm{d}^{3}x\sqrt{-g}J^{0},$ (15)
which can be interpreted as a particle number. This diagnostic is particularly useful for tracking the scalar-field content during the merger and for understanding the accretion dynamics onto the BH.
We also extract the outgoing GW radiation in terms of the Newman-Penrose scalar $\Psi_{4}$ Newman and Penrose (1962); Bishop and Rezzolla (2016); Radia et al. (2022) and decompose it into spin-weight s=2s=-2 spherical harmonics $Y^{-2}_{\ell m}$ Bruegmann et al. (2008) according to
$\Psi_{4}^{\ell m}(t,R_{\mathrm{ex}})=\int_{S^{2}}\Psi_{4}(t,R_{\rm ex},\theta,\phi)\overline{Y^{-2}_{\ell m}}(\theta,\phi)\sin\theta\mathrm{d}\theta\mathrm{d}\phi,$ (16)
where $S^{2}$ is a 2-sphere of fixed extraction radius $R_{\mathrm{ex}}$. By default, GRChombo performs this decomposition with respect to the grid’s $z$ axis as the polar axis. However, in our setup the binary undergoes a head-on collision strictly along the $x$ axis, so this choice obscures the underlying axisymmetry of the physical system. To recover the physically relevant axisymmetric modes aligned with the collision axis, we perform a passive rotation of the extracted multipole moments using the Wigner $D$-matrices:
${\Psi_{4}^{\prime}}^{\ell m^{\prime}}(t)=\sum_{m=-\ell}^{\ell}D^{l}_{m^{\prime}m}(\alpha,\beta,\gamma)\Psi_{4}^{\ell m}(t),$ (17)
where ${\Psi_{4}^{\prime}}^{\ell m^{\prime}}$ are the modes in the rotated frame and the Euler angles are $(\alpha,\beta,\gamma)=(0,\pi/2,0)$. In what follows, we drop the prime and refer to the rotated modes simply as $\Psi_{4}^{\ell m}$.
We also compute the cumulative GW energy $E_{\mathrm{GW}}$ from the Newman-Penrose scalar multipoles $\Psi_{4}^{\ell m}$ via
$\dot{E}_{\mathrm{GW}}^{\ell m}(t)$ $=\frac{R_{\mathrm{ex}}^{2}}{16\pi}\left|\int_{-\infty}^{t}\Psi_{4}^{\ell m}\mathrm{d}\tilde{t}\right|^{2},$ (18)
$\dot{E}_{\mathrm{GW}}(t)$ $=\sum_{\ell,m}\dot{E}_{\mathrm{GW}}^{\ell m}(t).$ (19)
We begin the time integration at $t=50$ to exclude any spurious or “junk” radiation present in the initial data.
III Initial data
III.1 Isolated boson star and black hole
To construct binary initial data, we first require stationary solutions for an isolated BS and an isolated BH in their respective rest frames. For convenience in converting to Cartesian coordinates and to avoid coordinate singularities, both compact objects are initially prepared in isotropic coordinates. The static, spherically symmetric metric takes the conformally flat form
$\mathrm{d}s^{2}=-\alpha^{2}(r)\mathrm{d}t^{2}+\psi^{4}(r)\delta_{ij}\mathrm{d}x^{i}\mathrm{d}x^{j},$ (20)
where $r=\sqrt{x^{2}+y^{2}+z^{2}}$ is the isotropic radial coordinate, $\alpha$ is the lapse function, and $\psi$ is the conformal factor. Since the shift vector vanishes in the rest frame ($\beta^{i}=0$), the extrinsic curvature $K_{ij}$ is identically zero.
For the boson star, the scalar field is assumed to be stationary and harmonic in time,
$\varphi(t,r)=A(r)e^{i\omega t},$ (21)
where $A(r)$ is a real radial amplitude and $\omega$ is a constant eigenfrequency. Substituting this ansatz and the isotropic metric into the Einstein-Klein-Gordon equations reduces the system to a set of coupled ordinary differential equations for $\alpha(r)$, $\psi(r)$, and $A(r)$. This system forms an eigenvalue problem, which we solve using a standard shooting method. By specifying the central scalar-field amplitude, $\varphi_{ctr}=A(0)$, the algorithm iteratively adjusts the frequency $\omega$ until the asymptotically flat boundary conditions $A\to 0$ and $\alpha,\psi\to 1$ as $r\to\infty$ are satisfied.
For the black hole, we use the standard isotropic Schwarzschild solution (puncture data) for a BH of bare mass $M_{\mathrm{BH}}$,
$\alpha_{BH}(r)$ $=\frac{1-M_{\mathrm{BH}}/2r}{1+M_{\mathrm{BH}}/2r},$ (22)
$\psi_{BH}(r)$ $=1+\frac{M_{\mathrm{BH}}}{2r},$
with the scalar field trivially satisfying $\varphi=0$ and the conjugate momentum $\Pi=0$. In the rest frame, the Arnowitt-Deser-Misner (ADM) Arnowitt et al. (2008) mass of the isolated BH is exactly $M_{\mathrm{BH}}$.
III.2 Lorentz boost
[…]
${\Lambda^{\tilde{\alpha}}}_{\mu}$ $=\begin{pmatrix}\cosh\eta&\sinh\eta&0&0\\ \sinh\eta&\cosh\eta&0&0\\ 0&0&1&0\\ 0&0&0&1\end{pmatrix}$ (23)
$\quad\Leftrightarrow\quad{\Lambda^{\mu}}_{\tilde{\alpha}}$ $=\begin{pmatrix}\cosh\eta&-\sinh\eta&0&0\\ -\sinh\eta&\cosh\eta&0&0\\ 0&0&1&0\\ 0&0&0&1\end{pmatrix}$.
Applying the coordinate transformation $\tilde{x}^{\tilde{\alpha}}={\Lambda^{\tilde{\alpha}}}_{\mu}x^{\mu}$ and the corresponding tensor transformation $\tilde{g}_{\tilde{\alpha}\tilde{\beta}}={\Lambda^{\mu}}_{\tilde{\alpha}}{\Lambda^{\nu}}_{\tilde{\beta}}g_{\mu\nu}$ to the isotropic metric, and evaluating the resulting fields on the initial hypersurface $\tilde{t}=0$, yields the exact $3+1$ ADM variables in the boosted frame. The boosted lapse $\tilde{\alpha}$ and shift vector $\tilde{\beta}^{\tilde{i}}$ are
$\tilde{\alpha}^{2}$ $=\frac{\psi^{4}\alpha^{2}}{\psi^{4}\cosh^{2}\eta-\alpha^{2}\sinh^{2}\eta},$ (24)
$\tilde{\beta}^{\tilde{x}}$ $=\frac{\sinh\eta\cosh\eta(\alpha^{2}-\psi^{4})}{\psi^{4}\cosh^{2}\eta-\alpha^{2}\sinh^{2}\eta},\quad\tilde{\beta}^{\tilde{y}}=\tilde{\beta}^{\tilde{z}}=0.$ (25)
The boosted spatial metric $\tilde{\gamma}_{ij}$ is no longer conformally flat and becomes
$\tilde{\gamma}_{\tilde{i}\tilde{j}}=\begin{pmatrix}\psi^{4}\cosh^{2}\eta-\alpha^{2}\sinh^{2}\eta&0&0\\ 0&\psi^{4}&0\\ 0&0&\psi^{4}\end{pmatrix}$. (26)
The extrinsic curvature are computed analytically from $\tilde{K}_{\tilde{i}\tilde{j}}=-\frac{1}{2\tilde{\alpha}}(\partial_{\tilde{t}}-\mathcal{L}_{\tilde{\beta}})\tilde{\gamma}_{\tilde{i}\tilde{j}}$ Croft (2022). The nonzero components are
$\tilde{K}_{\tilde{x}\tilde{x}}$ $=\tilde{\alpha}\sinh\eta\cosh^{2}\eta\frac{x}{r}\left(2\frac{\alpha^{\prime}}{\alpha}-2\frac{\psi^{\prime}}{\psi}-v^{2}\frac{\alpha\alpha^{\prime}}{\psi^{4}}\right),$ (27)
$\tilde{K}_{\tilde{x}\tilde{y}}$ $=\tilde{K}_{\tilde{y}\tilde{x}}=\tilde{\alpha}\sinh\eta\cosh\eta\frac{y}{r}\left(\frac{\alpha^{\prime}}{\alpha}-2\frac{\psi^{\prime}}{\psi}\right),$
$\tilde{K}_{\tilde{x}\tilde{z}}$ $=\tilde{K}_{\tilde{z}\tilde{x}}=\tilde{\alpha}\sinh\eta\cosh\eta\frac{z}{r}\left(\frac{\alpha^{\prime}}{\alpha}-2\frac{\psi^{\prime}}{\psi}\right),$
$\tilde{K}_{\tilde{y}\tilde{y}}$ $=\tilde{K}_{\tilde{z}\tilde{z}}=2\tilde{\alpha}\sinh\eta\frac{x}{r}\frac{\psi^{\prime}}{\psi},$
where $x=\tilde{x}\cosh\eta$, $y=\tilde{y}$, $z=\tilde{z}$, and $r=\sqrt{x^{2}+y^{2}+z^{2}}$. Primes denote derivatives with respect to $r$.
For the BS, the scalar field and its conjugate momentum must also be transformed to the boosted frame. Evaluated on the $\tilde{t}=0$ slice, the scalar field acquires a spatially modulated phase:
$\varphi(\tilde{\bm{x}})=\varphi(\bm{x})=A(r)e^{-i\omega\tilde{x}\sinh\eta}.$ (28)
[…]
$\quad\left.-\frac{\tilde{x}\cosh\eta}{\tilde{r}}A^{\prime}(\sinh\eta+\tilde{\beta}^{\tilde{x}}\cosh\eta)\right]e^{-i\omega\tilde{x}\sinh\eta}.$
This boost formalism is applied independently to the numerically integrated BS solution and the exact Schwarzschild BH solution (for which $\varphi=0$ and $\Pi=0$), preparing both objects for the subsequent superposition procedure. In the following, we focus on the lab-frame variables and drop the tildes for notational simplicity, with the understanding that all quantities now refer to the boosted frame.
III.3 Plain superposition and its malaise
[…]
$\gamma_{ij}$ $=\gamma_{ij}^{\mathrm{A}}+\gamma_{ij}^{\mathrm{B}}-\delta_{ij},$ (30a)
$K_{ij}$ $=\gamma_{m(i}\left[K_{j)n}^{\mathrm{A}}\gamma_{\mathrm{A}}^{mn}+K_{j)n}^{\mathrm{B}}\gamma_{\mathrm{B}}^{nm}\right],$ (30b)
$\varphi$ $=\varphi_{\mathrm{A}}+\varphi_{\mathrm{B}},$ (30c)
$\Pi$ $=\Pi_{\mathrm{A}}+\Pi_{\mathrm{B}}.$ (30d)
Here, the flat metric $\delta_{ij}$ is subtracted to ensure the correct asymptotic behavior at spatial infinity. For a BS-BH binary, we identify A with the BS and B with the BH, so that $\varphi_{\mathrm{B}}=\Pi_{\mathrm{B}}=0$. This construction is commonly referred to as the “plain superposition” method and has been widely used in numerical-relativity simulations of BH-BH Shibata et al. (2008); Okawa et al. (2011); Sperhake et al. (2019), BS-BS Palenzuela et al. (2007, 2008, 2017), and BS-BH Cardoso et al. (2022b); Zhong et al. (2023) systems. For pure BH spacetimes, this approximation has proved remarkably successful. However, because BSs are extended, horizonless objects supported by a delicate balance between scalar-field pressure and gravity, this simple construction can generate severe unphysical artifacts for them Helfer et al. (2019, 2022); Evstafyeva et al. (2023b). In particular, BSs are highly vulnerable to superposition artifacts near their centers because they lack the protective character of a BH horizon.
The main issue arises in the expression for the spatial metric $\gamma_{ij}$, Eq. (30a). At the center of the BS, $x^{i}=x^{i}_{\mathrm{A}}$, the gravitational field of the companion BH has not completely decayed to flat Minkowski space at finite initial separation $d$. As a result, the metric at the BS center is artificially perturbed by
$\delta\gamma_{ij}=\gamma_{ij}^{\mathrm{B}}(x^{i}_{\mathrm{A}})-\delta_{ij},$ (31)
away from the isolated equilibrium value $\gamma_{ij}^{\mathrm{A}}(x_{\mathrm{A}}^{i})$. Physically, this perturbation distorts the volume element $\sqrt{\gamma}$ at the BS core, effectively injecting spurious mass-energy and compressing the BS away from its delicate equilibrium. In practice, this artificial compression increases the Hamiltonian-constraint violation, induces unphysical radial oscillations, alters the GW signal, and, for sufficiently compact models (especially solitonic BSs), can trigger premature collapse into a BH before the physical merger begins. This is the principal “malaise” of plain superposition in BS-BH initial data: even when the asymptotic behavior is acceptable, the local BS core can already be driven away from equilibrium at $t=0$.
In Refs. Cardoso et al. (2022b); Zhong et al. (2023), the BH is much smaller than the BS and the initial separation is sufficiently large, so the superposition artifacts are relatively mild and do not significantly affect the dynamics. However, as we move toward more comparable mass ratios and smaller initial separations in order to systematically explore head-on collision dynamics, these artifacts become increasingly severe and can even induce premature collapse. It is therefore crucial to develop an improved superposition method that mitigates these artifacts and provides more accurate initial data for BS-BH binaries. This is the subject of the next subsection.
III.4 Improved superposition for boson star-black hole binaries
For the BS-BS case, Ref. Helfer et al. (2022) proposed an improved superposition method for equal-mass binaries that effectively removes the superposition artifacts by modifying the metric superposition to
$\gamma_{ij}=\gamma_{ij}^{\mathrm{A}}+\gamma_{ij}^{\mathrm{B}}-\gamma_{ij}^{\mathrm{B}}(x_{\mathrm{A}}^{i})=\gamma_{ij}^{\mathrm{A}}+\gamma_{ij}^{\mathrm{B}}-\gamma_{ij}^{\mathrm{A}}(x_{\mathrm{B}}^{i}).$ (32)
This method was later generalized to unequal-mass binaries in Ref. Evstafyeva et al. (2023b). The key idea is to retain the conformal metric of the plain superposition and modify only its conformal factor so that the BS center recovers the isolated-star volume element. The detailed prescription is given in Sec. 4.2 of Ref. Evstafyeva et al. (2023b); here we summarize the main steps.
Starting from the plainly superposed metric in Eq. (30a), we define a conformal factor $\lambda$ by
$\tilde{\gamma}_{ij}=\lambda^{-1}\gamma_{ij},\quad\lambda=\gamma^{1/3},$ (33)
where $\det\tilde{\gamma}_{ij}=1$ by construction, so that $\lambda=\chi^{-1}$. The improved method leaves $\tilde{\gamma}_{ij}$ unchanged and modifies only $\lambda$ to restore the equilibrium volume element at the object centers. For two generic objects A and B, the target conditions are
$\lambda_{\mathrm{new}}(x_{\mathrm{A}}^{i})$ $=\lambda(x_{\mathrm{A}}^{i})+\delta\lambda(x_{\mathrm{A}}^{i})=\lambda_{\mathrm{A}}(x_{\mathrm{A}}^{i}),$ (34)
$\lambda_{\mathrm{new}}(x_{\mathrm{B}}^{i})$ $=\lambda(x_{\mathrm{B}}^{i})+\delta\lambda(x_{\mathrm{B}}^{i})=\lambda_{\mathrm{B}}(x_{\mathrm{B}}^{i}).$ (35)
With weight functions $w_{\mathrm{A}}(x^{i})$ and $w_{\mathrm{B}}(x^{i})$, one introduces the ansatz
$\lambda_{\mathrm{new}}(x^{i})=\lambda(x^{i})+w_{\mathrm{A}}(x^{i})h_{\mathrm{A}}+w_{\mathrm{B}}(x^{i})h_{\mathrm{B}},$ (36)
where the coefficients $h_{\mathrm{A}}$ and $h_{\mathrm{B}}$ are determined by the resulting $2\times 2$ linear system,
$h_{\mathrm{A}}$ $=\frac{-w_{\mathrm{B}}(x_{\mathrm{A}}^{i})\delta\lambda(x_{\mathrm{B}}^{i})+w_{\mathrm{B}}(x_{\mathrm{B}}^{i})\delta\lambda(x_{\mathrm{A}}^{i})}{w_{\mathrm{A}}(x_{\mathrm{A}}^{i})w_{\mathrm{B}}(x_{\mathrm{B}}^{i})-w_{\mathrm{A}}(x_{\mathrm{B}}^{i})w_{\mathrm{B}}(x_{\mathrm{A}}^{i})}\,$, (37a)
$h_{\mathrm{B}}$ $=\frac{w_{\mathrm{A}}(x_{\mathrm{A}}^{i})\delta\lambda(x_{\mathrm{B}}^{i})-w_{\mathrm{A}}(x_{\mathrm{B}}^{i})\delta\lambda(x_{\mathrm{A}}^{i})}{w_{\mathrm{A}}(x_{\mathrm{A}}^{i})w_{\mathrm{B}}(x_{\mathrm{B}}^{i})-w_{\mathrm{A}}(x_{\mathrm{B}}^{i})w_{\mathrm{B}}(x_{\mathrm{A}}^{i})}.$ (37b)
The corrected metric is then reconstructed as
$\gamma^{\mathrm{new}}_{ij}=\left(\frac{\lambda_{\mathrm{new}}}{\lambda}\right)\gamma_{ij}=\frac{\lambda_{\mathrm{new}}}{\gamma^{1/3}}\gamma_{ij},$ (38)
and a convenient asymptotically flat choice for the weight functions is
[…]
where $R_{\mathrm{J}}$ is a tunable parameter that controls the spatial extent of the correction.
For our BS-BH problem, we adopt a similar strategy to remove the superposition artifacts at the BS center, with $\mathrm{A}\equiv\mathrm{BS}$ and $\mathrm{B}\equiv\mathrm{BH}$. However, applying this scheme to a heterogeneous BS-BH binary requires a careful treatment of the BH puncture. First, the conformal factor of the puncture BH diverges at its center, $\lambda_{\mathrm{BH}}(x_{\mathrm{BH}}^{i})=\psi_{\mathrm{BH}}^{4}(x_{\mathrm{BH}}^{i})\to\infty$ as $r\to 0$, so the relative error in the conformal volume element induced by the smooth BS background vanishes in the continuum limit. Thus, the required correction at the BH center is identically zero: $\delta\lambda_{\mathrm{BH}}=\lambda_{\mathrm{BH}}(x_{\mathrm{BH}}^{i})-\lambda(x_{\mathrm{BH}}^{i})\to 0$. Second, the BH interior is inherently immune to superposition artifacts because of the presence of the horizon. Therefore, we may choose the BH correction radius to satisfy $R_{\mathrm{BH}}\to 0$, which implies $w_{\mathrm{BH}}(x_{\mathrm{BH}}^{i})\to\infty$.
Taking these two exact physical limits, $\delta\lambda_{\mathrm{BH}}=0$ and $R_{\mathrm{BH}}\to 0$, within the full $2\times 2$ system, Eq. (37) reduces to a one-body correction centered on the BS,
$h_{\mathrm{BS}}=\frac{\delta\lambda_{\mathrm{BS}}(x_{\mathrm{BS}}^{i})}{w_{\mathrm{BS}}(x_{\mathrm{BS}}^{i})},\quad h_{\mathrm{BH}}=0,$ (40)
where $\delta\lambda_{\mathrm{BS}}=\lambda_{\mathrm{BS}}(x_{\mathrm{BS}}^{i})-\lambda(x_{\mathrm{BS}}^{i})$. Therefore,
$\lambda_{\mathrm{new}}(x^{i})=\lambda(x^{i})+w_{\mathrm{BS}}(x^{i})\frac{\delta\lambda_{\mathrm{BS}}(x_{\mathrm{BS}}^{i})}{w_{\mathrm{BS}}(x_{\mathrm{BS}}^{i})}.$ (41)
The improved spatial metric then follows from Eq. (38). This construction preserves the plain-superposition structure away from the BS core while enforcing the correct central volume element, which is the key ingredient for curing the spurious distortion discussed above. This BS-centered one-body conformal-factor correction is the BS-BH limit of the unequal-mass two-body BS-BS prescription, and in our simulations it significantly reduces the initial constraint violations and removes the spurious premature collapse seen in plain superposition, as we will demonstrate in the next section.
To complete the specification of the initial data for the $3+1$ evolution, we must also prescribe the initial gauge variables, even though their values do not affect the physical content of the initial spatial hypersurface. For the BS-BS case, one could similarly construct a superposition for the lapse $\alpha$ and shift vector $\beta^{i}$ as in Eq. (30). However, to stably evolve the BS-BH and BH-BH binaries in the moving-puncture framework, we instead initialize them using the standard pre-collapsed prescription:
$\alpha=\sqrt{\chi}=\sqrt{\lambda_{\mathrm{new}}^{-1}},\quad\beta^{i}=0.$ (42)
Although residual constraint violations remain in the initial data because of the superposition procedure, the improved method substantially reduces the unphysical artifacts relative to plain superposition. As shown in Ref. Atteneder et al. (2024), the dominant ingredient for improving the physical diagnostics, including the scalar-field dynamics and the extracted GW signal, is the volume-element correction at the BS center, while subsequent elliptic solving of the constraint equations provides only a subdominant improvement. In addition, the CCZ4 formulation used in our simulations includes built-in constraint damping, which further mitigates the impact of the residual initial violations on the subsequent evolution. We therefore leave the construction of fully constraint-solved initial data for future work and expect the improved superposition method to provide a substantial improvement over plain superposition in our BS-BH binary simulations.
IV Numerical results
Before presenting the results of our BS-BH binary simulations, we first summarize the initial-data configurations considered in this work. We focus on solitonic BSs with $\sigma_{0}=0.2$, which allows us to study relatively compact configurations. We consider three BS models with central amplitudes $A(0)=0.17$, $0.147$, and $0.1$, corresponding to highly compact, moderately compact, and low-compactness BSs, respectively. For simplicity, we label them H, M, and L. The corresponding properties of these BSs are listed in Table 1.
Model $A(0)$ $\mu M_{\mathrm{BS}}$ $\omega/\mu$ $\mu r_{99}$ $\mathcal{C}$
BS-170 0.17 0.7134 0.4392 3.3224 0.2146
BS-147 0.147 0.3609 0.6780 4.2557 0.0848
BS-100 0.1 0.2700 0.8506 6.1717 0.0437
Table 1: Solitonic BS models with $\sigma_{0}=0.2$ considered in this work. Here, $A(0)$ denotes the central scalar-field amplitude, $M_{\mathrm{BS}}$ the ADM mass of the BS, $\omega$ the frequency of the ground-state solution, $r_{99}$ the radius (in isotropic coordinates) containing $99\%$ of the BS mass, and $\mathcal{C}\equiv M_{\mathrm{BS}}/r_{99}$ is a measure of compactness. For this potential, the maximum BS mass is $\mu M_{\mathrm{BS}}=0.7212$.
For the BS-BH binaries, we consider mass ratios $q=M_{\mathrm{BS}}/M_{\mathrm{BH}}=0.38$, $0.5$, 11, 22, and $2.64$, and we also include several BS-BS and BH-BH cases for comparison. To test the validity of the improved superposition method, we additionally include a plain-superposition case for the equal-mass BS-BH binary and vary the initial separation to assess the severity of the artifacts. The parameters of the binary configurations are listed in Table 2. The initial velocities are chosen so that the total initial momentum vanishes, while the initial relative velocity is kept fixed at $\Delta v=v_{A}-v_{B}=0.2$ for all cases. For all BS-BH binaries, we use the improved superposition method described in Sec. III.4. For BSBH-q1-H, we also include the plain superposition method to illustrate the severity of the artifacts it produces. For the BS-BS binaries, we use the improved superposition method of Ref. Evstafyeva et al. (2023b) for all cases. We uniformly set the correction-radius parameter to $R_{\mathrm{BS}}=50$ for all BSs, which works well for the configurations considered here. For the BH-BH binaries, we use the plain superposition method, which is known to be sufficient for BH-BH initial data. For runs with $M_{\mathrm{BH}}=0.2700$, we use an additional refinement level to resolve the smaller BH.
Run $q$ Object A Object B $v_{x}^{\mathrm{A}}$ $v_{x}^{\mathrm{B}}$ $d$ $M$ Initial data
BSBH-q1-H 11 BS-170 $M_{\mathrm{BH}}=0.7134$ $0.1$ 0.1-0.1 $20,40,60,80$ $1.4340$ Improved/Plain
BSBH-q1-M 11 BS-147 $M_{\mathrm{BH}}=0.3609$ $0.1$ 0.1-0.1 $40$ $0.7254$ Improved
BSBH-q1-L 11 BS-100 $M_{\mathrm{BH}}=0.2700$ $0.1$ 0.1-0.1 $40$ $0.5427$ Improved
BSBH-q038 $0.38$ BS-100 $M_{\mathrm{BH}}=0.7134$ $0.1447$ 0.0553-0.0553 $40$ $0.9874$ Improved
BSBH-q05 $0.5$ BS-147 $M_{\mathrm{BH}}=0.7134$ $0.1325$ 0.0675-0.0675 $40$ $1.0791$ Improved
BSBH-q2 22 BS-170 $M_{\mathrm{BH}}=0.3609$ $0.0675$ 0.1325-0.1325 $40$ $1.0791$ Improved
BSBH-q2.64 $2.64$ BS-170 $M_{\mathrm{BH}}=0.2700$ $0.0553$ 0.1447-0.1447 $40$ $0.9874$ Improved
BSBS-q1-H 11 BS-170 BS-170 $0.1$ 0.1-0.1 $40$ $1.4340$ Improved
BSBS-q1-M 11 BS-147 BS-147 $0.1$ 0.1-0.1 $40$ $0.7254$ Improved
BSBS-q1-L 11 BS-100 BS-100 $0.1$ 0.1-0.1 $40$ $0.5427$ Improved
BSBS-q038 $0.38$ BS-100 BS-170 $0.1447$ 0.0553-0.0553 $40$ $0.9874$ Improved
BSBS-q05 $0.5$ BS-147 BS-170 $0.1325$ 0.0675-0.0675 $40$ $1.0791$ Improved
BHBH-q1 11 $M_{\mathrm{BH}}=0.7134$ $M_{\mathrm{BH}}=0.7134$ $0.1$ 0.1-0.1 $40$ $1.4340$ Plain
BHBH-q038 $0.38$ $M_{\mathrm{BH}}=0.2700$ $M_{\mathrm{BH}}=0.7134$ $0.1447$ 0.0553-0.0553 $40$ $0.9874$ Plain
BHBH-q05 $0.5$ $M_{\mathrm{BH}}=0.3609$ $M_{\mathrm{BH}}=0.7134$ $0.1325$ 0.0675-0.0675 $40$ $1.0791$ Plain
Table 2: Binary initial-data configurations considered in this work. Each run label begins with the binary type and ends with a suffix denoting the mass ratio $q=M_{\mathrm{A}}/M_{\mathrm{B}}$. Here, $v_{x}^{\mathrm{J}}$ for $\mathrm{J}\in\{\mathrm{A},\mathrm{B}\}$ denotes the initial boost velocity, with Lorentz factor $\gamma_{\mathrm{J}}=1/\sqrt{1-{v_{x}^{\mathrm{J}}}^{2}}$. The initial velocities are chosen so that the total initial momentum $P_{x}=M_{\mathrm{A}}\gamma_{\mathrm{A}}v_{x}^{\mathrm{A}}+M_{\mathrm{B}}\gamma_{\mathrm{B}}v_{x}^{\mathrm{B}}$ vanishes, while the initial relative velocity is fixed at $\Delta v=v_{x}^{\mathrm{A}}-v_{x}^{\mathrm{B}}=0.2$ for all runs. The total mass of the binary is denoted by $M=\gamma_{\mathrm{A}}M_{\mathrm{A}}+\gamma_{\mathrm{B}}M_{\mathrm{B}}$. For the BS-BH binaries, we use the improved superposition method described in Sec. III.4 for all runs except BSBH-q1-H, for which we also include the plain superposition method to demonstrate its artifacts. For the BS-BS binaries, we use the improved superposition method of Ref. Evstafyeva et al. (2023b) for all runs. For the BH-BH binaries, we use the plain superposition method for all runs, since it does not suffer from the BS-specific superposition artifacts.
IV.1 Comparison of initial data
IV.1.1 Initial constraint violations
Figure 1: Comparison of the initial Hamiltonian (top panel) and momentum (bottom panel) constraint violations along the collision axis for the configuration BSBH-q1-H with $d=20$, generated using the plain (solid lines) and improved (dashed lines) superposition methods, respectively. The improved superposition method significantly reduces the Hamiltonian-constraint violation in the vicinity of both the BS (at $x=246$) and the BH (at $x=266$) relative to the plain superposition method.
Since the plain superposition method can distort the volume element at the BS center and thereby inject spurious mass-energy, one expects the improved superposition method to reduce the Hamiltonian-constraint violation significantly. To demonstrate this, we plot the initial Hamiltonian-constraint violation along the collision axis for the configuration BSBH-q1-H with $d=20$ in the top panel of Fig. 1, comparing the plain and improved superposition methods with solid and dashed lines, respectively. We find that the improved superposition method reduces the Hamiltonian-constraint violation at the BS center (at $x=246$) by several orders of magnitude relative to the plain superposition method. Moreover, the Hamiltonian-constraint violation is also substantially reduced in the vicinity of the BH puncture (at $x=266$), even though no correction is applied there (unlike the TwoPunctures correction used in Ref. Marks et al. (2026)). The momentum-constraint violation, shown in the bottom panel of Fig. 1, is also mildly reduced near the BS. These results demonstrate the effectiveness of the improved superposition method in reducing constraint violations and providing more accurate initial data for BS-BH binaries.
IV.1.2 Evolution of the scalar field and premature collapse
We now compare the time evolution of the scalar field for the configuration BSBH-q1-H generated using the plain and improved superposition methods. The model BS-170 is a highly compact BS, lying below but close to the maximum-mass threshold and therefore close to the instability threshold. Fig. 2 shows the time evolution of the central scalar-field amplitude $|\varphi_{c}(t)|$ for the plain (solid lines) and improved (dashed lines) superposition methods at different initial separations. For the plain-superposition runs with $d=20$, $40$, and $60$, the scalar amplitude increases significantly from its initial value and then rapidly drops to a near-zero level, indicating premature collapse of the BS into a BH before the physical merger dynamics begins. The formation of a BH is also confirmed by the horizon finder. For the largest initial separation, $d=80$, premature collapse is avoided, but the scalar amplitude still exhibits significant unphysical radial oscillations, with the equilibrium value acting as the minimum value. In all cases, the improved superposition method removes the unphysical oscillations and prevents premature collapse, keeping the scalar amplitude close to its initial equilibrium value until the physical merger dynamics begins. These results demonstrate that BS-BH collisions exhibit the same type of spurious behavior caused by plain superposition as in BS-BS collisions, and that the improved superposition method effectively mitigates these artifacts.
Figure 2: Time evolution of the central scalar-field amplitude $|\varphi_{c}(t)|$ for the configuration BSBH-q1-H, generated using the plain (solid lines) and improved (dashed lines) superposition methods at different initial separations. The plain superposition method leads to strong unphysical radial oscillations and, for smaller initial separations, premature collapse of the BS into a BH. The improved superposition method removes these artifacts and keeps the scalar amplitude close to its initial equilibrium value until the physical merger dynamics begins.
IV.1.3 Radiated gravitational waveforms and energy
We now present the GW waveforms and radiated energy for the configuration BSBH-q1-H, comparing the plain and improved superposition methods at different initial separations. Because the plain superposition method injects spurious mass-energy into the BS and even triggers premature collapse at smaller separations, it is expected to produce more GW energy than the improved superposition method, which provides more accurate initial data. This is confirmed by the results shown in the top panel of Fig. 3. The increase in the radiated energy with increasing initial separation is physical and is associated with the larger collision velocity at merger due to the longer free-fall time, as discussed in Ref. Evstafyeva et al. (2023b) for the BS-BS case. However, the plain superposition method overestimates the radiated energy relative to the improved method, especially for smaller initial separations. As the initial separation increases, the difference between the two methods becomes less pronounced, and for $d=80$ both approaches yield compatible results because the BH gravitational tail has sufficiently decayed at the BS center.
Figure 3: Comparison of the radiated GW energy (top panel) and the $(2,0)$ mode of the Newman-Penrose scalar $\Psi_{4}$ (bottom panel) for the configuration BSBH-q1-H, generated using the plain (solid lines) and improved (dashed lines) superposition methods at different initial separations. Relative to the improved method, the plain superposition method overestimates the radiated energy and introduces unphysical distortions in the waveform, especially at smaller initial separations. As the initial separation increases, the two methods yield more compatible results.
In the bottom panel of Fig. 3, we show the $(2,0)$ mode of the Newman-Penrose scalar $\Psi_{4}$ for the same configurations. The amplitude of the $(2,0)$ mode is slightly higher for the improved superposition method than for the plain superposition method, although the difference is not large. However, the waveforms produced by the plain superposition method are significantly stretched in time relative to those from the improved method, which leads to a larger radiated energy after time integration. This stretching effect is more pronounced at smaller initial separations, whereas for larger separations the waveforms from the two methods become more similar. These results show that the plain superposition method not only overestimates the radiated energy but also introduces unphysical distortions into the waveform. The improved superposition method is therefore essential for obtaining accurate GW signals and energy estimates from BS-BH binary simulations, particularly when the initial separation is small and the superposition artifacts are strongest.
IV.2 Equal-mass BS-BH collisions
After demonstrating the effectiveness of the improved superposition method in mitigating artifacts and obtaining more accurate GWs for BS-BH binaries, we now present the results of our BS-BH simulations. We first focus on the equal-mass case with $q=1$ and compare the gravitational waveforms for BS-BH, BS-BS, and BH-BH collisions across different BS models.
The top panel of Fig. 4 shows the radiated GW energy for equal-mass BS-BH, BS-BS, and BH-BH collisions with an initial separation of $d=40$. For the BS-BH collisions (solid lines), our results agree with the recent findings of Marks et al. Marks et al. (2026): the GW radiation efficiency increases with the compactness of the BS, and the most compact BS-BH system approaches the BH-BH limit from below (see the BS-170 model). By contrast, for the BS-BS collisions (dashed lines) with the same BS models, a strikingly different trend emerges: in both cases, the radiated energy exceeds that of the BH-BH case and decreases as the compactness increases. Moreover, the radiated energy for the medium- and low-compactness BS-BS collisions is much larger than that of the BH-BH case. This behavior reflects the fact that GW emission efficiency is determined not only by compactness, but also by the tidal deformability Sennett et al. (2017) of the compact objects. For a detailed discussion of GW emission from BS-BS binaries, we refer the reader to Refs. Ge (2024); Ge et al. (2025); Ge (2025). In contrast, in BS-BH collisions the BH rapidly accretes the extended scalar cloud of the BS, thereby suppressing the kinetic energy that would otherwise be radiated as GWs. Therefore, one should not naively expect the GWs emitted by a BS-BH binary to lie between those emitted by the corresponding BS-BS and BH-BH systems.
Figure 4: Comparison of the radiated GW energy (top panel) and the $(2,0)$ (middle panel) and $(3,0)$ (bottom panel) modes of $\Psi_{4}$ for equal-mass BS-BH, BS-BS, and BH-BH collisions with different BS models. The BS-BH collisions exhibit a unique excitation of the $(3,0)$ mode due to the asymmetry between the BS and the BH, which is absent in both the BS-BS and BH-BH collisions.
The middle panel of Fig. 4 shows the $(2,0)$ mode of the Newman-Penrose scalar $\Psi_{4}$ for the same configurations. All three binary types exhibit a similar qualitative structure—a primary merger pulse followed by a ringdown phase—but they differ markedly in their peak amplitudes and pulse widths. For the highly compact case (H), the BS-BH and BS-BS waveforms are morphologically almost indistinguishable from the BH-BH baseline, reinforcing the “BH mimicker” paradigm in the highly compact regime. However, for the less compact configurations (M and L), a pronounced contrast appears. The BS-BS collisions exhibit a strongly amplified and sharply peaked primary pulse, whereas the amplitudes of the corresponding BS-BH waveforms are substantially suppressed.
Perhaps the most striking distinguishing feature of mixed BS-BH binaries is the excitation of higher-order modes. In the bottom panel of Fig. 4, we plot the $(3,0)$ multipole of the Newman-Penrose scalar. For equal-mass BH-BH and BS-BS binaries, the $(3,0)$ mode vanishes identically, as required by reflection symmetry along the collision axis. Remarkably, for BS-BH binaries, even when the macroscopic masses are equal, the intrinsic asymmetry of the system—a singular horizon interacting with a diffuse scalar cloud—breaks reflection symmetry and excites the $(3,0)$ mode. Our results show that, for equal-mass BS-BH collisions, the amplitude of the $(3,0)$ mode is about $5\%$$10\%$ of that of the $(2,0)$ mode and increases as the compactness of the BS decreases, because a less compact BS makes the binary more asymmetric. This distinctive excitation in equal-mass mixed binaries provides a promising “smoking-gun” signature for breaking the observational degeneracy between BHs and BSs. It also highlights the importance of studying heterogeneous binaries such as BS-BH systems, which can exhibit waveform features absent in homogeneous binaries.
We also plot the time evolution of the Noether charge $N$ for the equal-mass BS-BH and BS-BS collisions in Fig. 5. For the BS-BH collisions, the Noether charge nearly vanishes after the merger, indicating that the BS is almost completely disrupted and accreted by the BH, with only a small fraction ($10\%$) of the scalar field remaining outside the horizon. Moreover, the amount of remnant scalar field increases as the compactness of the BS decreases, consistent with Refs. Clough et al. (2018); Marks et al. (2026). For the BS-BS collisions, the Noether charge also nearly vanishes after the merger when the final state is a BH (for the high- and medium-compactness cases), whereas for the low-compactness case the final state is a BS, and the Noether charge remains constant after the merger.
Figure 5: Time evolution of the Noether charge $N$ for equal-mass BS-BH (solid lines) and BS-BS (dashed lines) collisions with different BS models. For the BS-BH collisions, the remnant Noether charge after the merger increases as the compactness of the BS decreases.
IV.3 Unequal-mass BS-BH collisions
Figure 6: Comparison of the radiated GW energy (top panels), the $(2,0)$ mode (middle panels), and the $(3,0)$ mode (bottom panels) of $\Psi_{4}$ for unequal-mass BS-BH (solid lines) and BH-BH (dotted lines) collisions with mass ratios $q\in\{0.5,2\}$ (left column) and $q\in\{0.38,2.64\}$ (right column). Note that the $(3,0)$ modes for configurations with $q>1$ have been multiplied by 1-1 so that their initial polarities match those with $q<1$, for better visualization.
Before examining the GW emission from unequal-mass binaries, we must address a major distinction between BS-BS and BS-BH systems in their initial parameter space. As emphasized by Evstafyeva et al. Evstafyeva et al. (2023b), the merger dynamics and radiated energy of unequal-mass BS-BS binaries are highly sensitive to the relative scalar phase at the moment of collision. Because unequal-mass BSs have different characteristic frequencies ($\omega_{A}\neq\omega_{B}$), they inevitably accumulate a dynamical phase difference $\Delta\phi\approx(\omega_{A}-\omega_{B})t$ during the infall. Consequently, setting the initial phase difference to zero at $t=0$ leads to a merger phase—and therefore to constructive or destructive interference patterns—that depends strongly on the initial separation $d$. This strong phase degeneracy makes it difficult to define a representative GW energy for unequal-mass BS-BS collisions without an exhaustive scan over the phase space.
By contrast, the BH in a BS-BH binary carries no scalar phase. The interaction and accretion dynamics are governed solely by the BS stress-energy tensor, which is invariant under global $U(1)$ phase transformations. The merger dynamics of BS-BH systems are therefore intrinsically deterministic, robust, and independent of any phase offset. To ensure a rigorous and unambiguous comparison across binary parameters, we exclude the strongly phase-dependent BS-BS configurations from the following analysis of unequal-mass systems. Instead, we focus on contrasting the phase-independent BS-BH collisions with their pure BH-BH counterparts across different mass ratios $q$.
We first examine the radiated GW energy, shown in the top panels of Fig. 6 for mass ratios $q\in\{0.5,2\}$ (left) and $q\in\{0.38,2.64\}$ (right). When the BH is the heavier companion ($q<1$, orange solid lines), the radiative efficiency of the BS-BH binary is consistently lower than that of the corresponding BH-BH system (blue dotted lines), as in the equal-mass case. However, a more intricate nonlinear competition emerges when the highly compact BS is the heavier object ($q>1$, green solid lines). For $q=2$, the radiated energy of the BS-BH system slightly exceeds that of the pure BH-BH collision. This is consistent with the recent findings of Marks et al. Marks et al. (2026), who observed that sufficiently compact BS-BH collisions can radiate more efficiently than their BH-BH counterparts in the $q>1$ regime. Interestingly, when the mass ratio is increased further to $q=2.64$, the BS-BH energy drops slightly below the BH-BH baseline again. This behavior appears to contrast with Ref. Marks et al. (2026), where the positive excess in radiated energy for $q>1$ was found to grow with $q$ up to $q=3$. We attribute this difference to the distinct kinematic setups. Marks et al. adopted a constant initial velocity ($v=0.1$) for all objects, thereby assigning an artificially large kinetic energy to the more massive BS. In contrast, we use a rigorously controlled center-of-momentum frame ($P_{x}=0$, $\Delta v=0.2$), which is more physically realistic. In the $q=2.64$ configuration, the much lighter BH must move with a relatively large initial velocity ($v_{\mathrm{BH}}=0.1447$) to plunge into the massive, slowly moving BS ($v_{\mathrm{BS}}=0.0553$). This “bullet-like” piercing induces strong dynamical friction, transferring a substantial fraction of the BH kinetic energy into internal scalar excitations and scalar wakes rather than into quadrupolar GWs. Thus, in $q>1$ mixed binaries, the radiated GW energy reflects a delicate competition between the increased emission associated with the larger mass ratio and the dissipative effects of kinematic piercing, which suppress the radiation.
Despite these differences in the total radiated energy, the dominant $(2,0)$ quadrupolar waveforms (middle panels of Fig. 6) remain morphologically very similar across all configurations. Their temporal width, merger peak, and ringdown frequencies closely track those of the pure BH-BH baseline. This reaffirms the paradigm that, in the dominant GW channel, highly compact BSs are excellent BH mimickers, even in the unequal-mass regime.
Although the $(2,0)$ mode exhibits strong degeneracies, the higher-order $(3,0)$ multipole—which is driven by the asymmetry of the mass distribution—reveals much more pronounced morphological differences (bottom panels of Fig. 6). Unlike the equal-mass case, the $q\neq 1$ configurations break reflection symmetry and thus excite the $(3,0)$ mode in all systems, including the pure BH-BH baseline. A closer inspection of the $(3,0)$ multipole reveals a clear physical distinction between the $q>1$ and $q<1$ regimes, underscoring the central role of tidal disruption. When the highly compact BS is the heavier object ($q>1$), the BS-BH waveform is almost indistinguishable from the pure BH-BH baseline. This near-perfect agreement reflects the structural properties of these specific configurations: in both cases, the massive primary object is the highly compact BS-170. Its deep gravitational potential well and dense core make it robust against the tidal field of the plunging light BH. As a result, the piercing dynamics closely resemble the point-mass evolution of a BH-BH collision.
By contrast, a pronounced morphological deviation appears when the BH is the heavier object ($q<1$). Relative to the BH-BH baseline, the BS-BH $(3,0)$ waveforms exhibit a clear phase advance and amplitude enhancement, especially for $q=0.5$. This behavior can be directly linked to the microphysical accretion dynamics, as seen in the scalar-field modulus cross-sections in Fig. 7. In the $q=0.5$ case, the BS has already been compressed by the heavier BH at $t=145$. The BH nearly engulfs the BS by $t=150$, and a distinct elongated scalar wake forms on the trailing side of the horizon by $t=160$. This persistent, highly asymmetric mass current provides a large lever arm for the octupole moment and therefore amplifies the $(3,0)$ emission pulse. In contrast, when the light BH pierces the massive BS ($q=2$), the BH quickly accretes the BS without producing a significant trailing wake. The absence of such strong asymmetries explains why the $q=2$ configuration emits octupolar radiation that almost perfectly mimics the pure BH-BH baseline. At very late times ($t=200$), the residual scalar fields—the nascent “gravitational atoms”—display markedly different morphologies that encode their merger histories. For $q=2$, the scalar remnant settles into a highly left-right symmetric, dual-lobe configuration, whereas for $q=0.5$ the remnant remains strongly asymmetric, retaining a complex multipolar structure and a persistent bright one-sided tail.
Figure 7: Cross-sectional snapshots of the scalar-field modulus $|\varphi|$ in the $x$-$y$ plane for unequal-mass BS-BH collisions with mass ratios $q=0.5$ (top row) and $q=2$ (bottom row) at different times. The red circle marks the contour of constant conformal factor $\chi=1/16$, which corresponds to the value at the event horizon for a Schwarzschild BH in isotropic coordinates and serves as a rough estimate of the horizon location.
The amplitude enhancement in the $q=0.5$ case is visibly more pronounced than in the more extreme $q=0.38$ case, illustrating the competition between tidal-field strength, controlled by the mass ratio, and the structural integrity, controlled by the compactness, of the BS. A systematic study of this effect would require many additional simulations and is left for future work. These results demonstrate that higher-order modes provide a highly promising, mass-ratio-dependent observational signature for breaking waveform degeneracies between mixed BS-BH binaries and pure BH systems.
V Conclusions and discussion
In this work, we have presented a systematic numerical-relativity study of comparable-mass BS-BH head-on collisions, with emphasis on both initial-data construction and GW phenomenology. Our main technical advance is an improved BS-centered one-body conformal-factor correction for BS-BH binaries, obtained as the BS-BH limit of the unequal-mass BS-BS prescription. This construction is designed to restore the BS core volume element while preserving the plain-superposition structure away from the star.
Our principal findings can be summarized as follows.
First, at the level of the initial data, the improved prescription significantly suppresses the Hamiltonian-constraint violation near the BS core, and it also reduces the violation around the BH puncture, compared with plain superposition. More importantly, it removes the spurious early collapse and large unphysical oscillations of the central scalar amplitude that appear in plain-superposition evolutions for compact stars at modest separations. The resulting GW signals are correspondingly cleaner and more physically robust.
Second, for equal-mass binaries, BS-BH collisions show a clear compactness trend: the GW radiation efficiency increases with BS compactness and approaches the BH-BH limit for highly compact stars. In contrast, the equal-mass BS-BS collisions in our sample can radiate more strongly than the BH-BH case and follow a different compactness trend, highlighting that compactness alone does not control the emission efficiency across different binary classes. We also find that the $(3,0)$ mode provides a distinctive signature of mixed BS-BH systems: it is nonzero even in the equal-mass case due to the intrinsic asymmetry between a horizon and a scalar condensate, whereas it vanishes for equal-mass BH-BH and BS-BS head-on collisions.
Third, for unequal-mass BS-BH binaries, the dominant $(2,0)$ mode remains close to the BH-BH morphology in many cases, but the subdominant $(3,0)$ mode carries substantial discriminatory power. In particular, when the BH is the heavier object ($q<1$), stronger tidal disruption and asymmetric scalar wakes lead to earlier and larger octupolar emission. For $q>1$, the waveforms are closer to those of BH-BH collisions, and the total radiated energy reflects a competition between enhancement driven by the mass ratio and dissipation associated with piercing and accretion dynamics. This helps explain the non-monotonic behavior we observe across mass ratios.
From an observational perspective, these results reinforce a two-channel picture: (i) high compactness can make BS-BH systems partially degenerate with BH-BH systems in the dominant quadrupolar emission, but (ii) higher-order multipoles provide a key handle for breaking that degeneracy. In this sense, accurate initial data are not merely a numerical detail, but a prerequisite for extracting physically reliable multimode GW signatures.
Our study also has clear limitations. We have focused on head-on collisions, a single scalar-potential family, and a finite set of mass ratios and compactnesses. Extending this program to quasi-circular inspirals, broader potential models, including mini-BS baselines, and a larger parameter-space coverage is a natural next step. It will also be important to quantificationally compare against fully constraint-solved initial data and different initial-data correction prescriptions, such as the method of Ref. Marks et al. (2026).
Overall, our results show that comparable-mass BS-BH binaries constitute a rich and distinct class of compact-binary sources: they can mimic BH-BH behavior in some channels while exhibiting unique scalar-matter and higher-mode signatures in others. This dual character makes them a particularly promising target for future multimode GW analyses and for tests of the existence of bosonic compact objects in strong gravity.
Note added. While we were finalizing this work, Ref. Marks et al. (2026) appeared on arXiv, which also studied comparable-mass BS-BH binaries using numerical-relativity simulations. Their study and ours overlap in the broad goal of characterizing GW emission from BS-BH mergers, but they are complementary in methodology and diagnostics. First, we employ a BS-centered one-body conformal-factor correction derived from the unequal-mass BS-BS prescription, whereas Ref. Marks et al. (2026) uses a different initial-data treatment that combines a BS metric correction with a BH TwoPunctures correction. Second, we cross-compare the GW emission from BS-BH binaries with both BS-BS and BH-BH counterparts at matched kinematics, whereas Ref. Marks et al. (2026) focuses on the comparison between BS-BH and BH-BH binaries. In the overlapping regimes, both studies indicate that compactness is a key control parameter for GW emission, and that highly compact BS-BH systems can approach BH-BH-like GW signatures. Quantitative differences in some unequal-mass cases are plausibly attributable to different initial-data prescriptions and kinematic setups, in particular the choice of initial velocities and center-of-momentum conditions. Third, Ref. Marks et al. (2026) mainly focuses on the radiated GW energy and the dominant $(2,0)$ mode, while we also analyze the higher-order $(3,0)$ mode and identify it as a promising smoking-gun signature for breaking the degeneracy between BS-BH and BH-BH binaries. Overall, both studies provide valuable insights into GW emission from BS-BH binaries and highlight the importance of accurate initial data and comprehensive diagnostics for understanding these complex systems.
Acknowledgements.
This work is supported in part by the National Key Research and Development Program of China Grant No. 2021YFC2203004, No. 2021YFA0718304, and No. 2020YFC2201501, the National Natural Science Foundation of China Grants No. 12422502, No. 12547110, No. 12588101, No. 12235019, and No. 12447101, and the Science Research Grants from the China Manned Space Project with No. CMS-CSST-2025-A01.
Appendix A Convergence tests
We use the run BSBH-q1-H with $d=40$ to perform convergence tests of our numerical simulations. We use three different resolutions with finest grid spacings of $\Delta x_{l}=1/16$, $\Delta x_{m}=1/32$, and $\Delta x_{h}=1/64$, which we refer to as low, medium, and high resolutions, respectively. The convergence order is estimated by comparing the differences between results at different resolutions. The $p$th-order convergence factor is defined as $Q_{p}\equiv(\Delta x_{m}^{p}-\Delta x_{l}^{p})/(\Delta x_{h}^{p}-\Delta x_{m}^{p})$. In Fig. 8, we show the GW-energy differences between the different resolutions as a function of time. The differences between the high- and medium-resolution results are multiplied by the convergence factor $Q_{p}$ for $p=2$ and $p=3$. Comparing these with the differences between the low- and medium-resolution results, we find that the convergence order lies between 2 and 3. In Fig. 9, we show the $L_{2}$ norm of the Hamiltonian constraint violation (left panel) and the magnitude of the momentum-constraint violation (right panel) as functions of time for the three resolutions. We find that the constraint violations decrease with increasing resolution, further confirming the convergence of our simulations.
Figure 8: Convergence test of the radiated GW energy for the run BSBH-q1-H with $d=40$ at three different resolutions. The differences between the high- and medium-resolution results are multiplied by the convergence factor $Q_{p}$ for $p=2$ and $p=3$. The convergence order is estimated to lie between 2 and 3.
Figure 9: $L_{2}$ norm of the Hamiltonian constraint violation (left panel) and the magnitude of the momentum-constraint violation (right panel) as functions of time for the run BSBH-q1-H with $d=40$ at three different resolutions.